OpenStudy (anonymous):

differentiate: j(x) = x^2.4+e^2.4

OpenStudy (anonymous):

@ikram002p

OpenStudy (anonymous):

just bring down the power and subtract with 1?

OpenStudy (anonymous):

e is the math e

OpenStudy (ikram002p):

so like this \( j(x) = x^2.4+e^2.4 \) yep ur correct :D

OpenStudy (anonymous):

no its x^(2.4) + e^(2.4)

OpenStudy (anonymous):

so 2.4x^(1.4) +2.4e^(1.4)?

OpenStudy (ikram002p):

no lol e^2.4 is a nummer=constant

OpenStudy (ikram002p):

i thought both of them sry

OpenStudy (imstuck):

It would be 2.4x^1.4 + e^2.4. the derivative of e to the anything stays the same.

OpenStudy (ikram002p):

so for x part ur correct but for e^1.4=0

OpenStudy (anonymous):

oh thank you!

OpenStudy (ikram002p):

lol whats ur answer ?

OpenStudy (anonymous):

2.4x^(2.4-1) +e^2.4 2.4x^(1.4)+e^2.4

OpenStudy (ikram002p):

incorrect

OpenStudy (ikram002p):

e^2.4 is constant like C

OpenStudy (anonymous):

so 0 for e

OpenStudy (ikram002p):

yep

OpenStudy (anonymous):

2.4x^1.4 thats the answer

OpenStudy (ikram002p):

yess :)

OpenStudy (anonymous):

Thank you!!

OpenStudy (ikram002p):

np :)

OpenStudy (anonymous):

another question i will tag you

OpenStudy (ikram002p):

ok post